# Representation Learning and Identity Adversarial Training for Facial Behavior Understanding

### v1

@@ Abstract
Facial Action Unit (AU) detection has gained significant research attention as AUs contain complex expression information. In this paper, we unpack two fundamental factors in AU detection: data and subject identity regularization, respectively. Motivated by recent advances in foundation models, we highlight the importance of data and collect a diverse dataset Face9M, comprising 9 million facial images, from multiple public resources. Pretraining a masked autoencoder on Face9M yields strong performance in AU detection and facial expression tasks. We then show that subject identity in AU datasets provides a shortcut learning for the model and leads to sub-optimal solutions to AU predictions. To tackle this generic issue of AU tasks, we propose Identity Adversarial Training (IAT) and demonstrate that a strong IAT regularization is necessary to learn identity-invariant features. Furthermore, we elucidate the design space of IAT and empirically show that IAT circumvents the identity shortcut learning and results in a better solution. Our proposed methods, Facial Masked Autoencoder (FMAE) and IAT, are simple, generic and effective. Remarkably, the proposed FMAE-IAT approach achieves new state-of-the-art F1 scores on BP4D (67.1%), BP4D+ (66.8%), and DISFA (70.1%) databases, significantly outperforming previous work. We release the code and model at https://github.com/forever208/FMAE-IAT, the first open-sourced facial model pretrained on 9 million diverse images.
@@ 1 Introduction
The Facial Action Coding System (FACS) was proposed by Ekman for objectively encoding facial behavior through specific movements of facial muscles, named Action Units (AU) [19]. Compared with facial expression recognition (FER) [37, 84, 35] and valence and arousal estimation [85, 52, 55], action units offer a more nuanced and detailed understanding of human facial behavior, because multiple AUs can occur simultaneously at varying intensities, whereas FER typically involves single-label classification.
@@
This problem attracted considerable interest within the deep learning community [13, 32, 73, 57, 62, 74]. Many works used a facial region prior [39, 57, 13], introduced extra modalities [80, 70, 81], or incorporated the inherent AU relationships [38, 47, 76] to solve the AU detection task and achieved significant advancements. Diverging from these approaches, which often necessitate complex model designs or depend heavily on prior AU knowledge, in this paper, we focus on two fundamental factors that significantly contribute to the AU detection task: data and subject identity regularization, respectively.
@@
Recently, data has become pivotal in training foundation models [58, 2, 42, 65] and large language models [8, 1, 66, 15]. Following this trend, we introduce a large-scale facial dataset Face9M, curated and refined from publicly available datasets, for pretraining. Since most facial tasks require a fine-grained understanding of the face, we propose to do representation learning using Masked Autoencoders (MAE) [27] as masked pretraining results in lower level semantics than contrastive learning [5]. Our large-scale facial representation learning approach demonstrates excellent generalization and scalability in downstream tasks. Notably, our proposed Facial Masked Autoencoder (FMAE), pretrained on Face9M, sets new state-of-the-art benchmarks in both AU detection and FER tasks.
@@
Similar to the importance of data, domain knowledge and task-prior knowledge can be incorporated into the model in the form of regularization [26, 31, 53, 56] to improve task performance. Our key observation is that popular AU detection benchmarks (i.e. BP4D [83], BP4D+ [86], DISFA [51]) have at most 140 human subjects and 192,000 images, which means that each subject has hundreds of annotated images. This abundance can lead models to prefer simple, easily recognizable patterns over more complex but generalizable ones, as suggested by the shortcut learning theory [23, 29]. Therefore, we hypothesize that AU detection models tend to learn the subject identity features to infer the AUs, resulting in learning a trivial solution that does not generalize well. To verify our hypothesis, we employed the linear probing technique — adding a learnable linear layer to a trained AU model while freezing the network backbone —to measure identity recognition accuracy. The high accuracy (83%) we obtained clearly shows that the models effectively ‘memorize’ subject identities. To counteract the learning of identity-based features, we propose in this paper Identity Adversarial Training (IAT) for AU tasks by adding a linear identity prediction head and unlearning the identity feature using gradient reverse [20]. Further analysis shows that IAT significantly reduces the identity accuracy of linear probing and leads to better learning dynamics that avoid convergence to trivial solutions. This method further improves our AU models beyond the advantages brought by pretraining with a large-scale dataset.
@@
  • We illustrate the identity learning issue and propose the IAT method for AU detection.
  • We release the code and checkpoint of FMAE with various model sizes (small, base, large), hoping to facilitate all facial tasks.
@@ 3.1 Large-scale Facial MAE Pretraining
While the machine learning community has long established the importance of having rich and diverse data for training, recent successes in foundation models and large language models illustrated the full potential of pertaining [58, 42, 8, 1, 66]. In line with this, our research pivots towards a nuanced exploration of data diversity and quantification in the context of facial representation learning. This approach marks a departure from previous work that predominantly concentrated on refining representation learning algorithms. By prioritizing the scope and scale of data employed for facial pretraining, we aim to uncover more robust and universally applicable facial representations.
We first collect facial images from CelebA  [44], FFHQ  [33], VGGFace2 [11], CASIA-WebFace [77], MegaFace [34], EDFace-Celeb-1M [79], UMDFaces [6] and LAION-Face [89] datasets, because these datasets contain a massive number of identities collected in diverse scenarios. For instance, the facial images in UMDFaces also capture the upper body with various image sizes, while some datasets (FFHQ, CASIA-WebFace) mainly feature the center face. We then discard images whose width-height-ratio or height-width-ratio is larger than 1.5. Finally, the remaining images are resized to 224*224. The whole process yields 9 million facial images (termed Face9M) which will be used for self-supervised facial pretraining.
Regarding representation learning methods, We apply Masked Image Modeling [27] as it tends to learn more fine-grained features [5] than contrastive learning, which benefits facial behavior understanding. Specifically, we utilize Face9M to train a masked autoencoder (MAE) by the mean squared error between the reconstructed and original images in the pixel space. The decoder of MAE is then discarded for downstream tasks.
@@ 3.2 Identity Adversarial Training
One of our key findings in this paper is that, the limited number of subjects in AU datasets makes identity recognition a trivial task and provides a shortcut learning path, resulting in a AU model that contains identity-related features and does not generalize well (see Sec. 5). Motivated by the gradient reverse in domain adaption [20], we propose to apply gradient reverse on AU detection to learn identity-invariant features, aiming at better model generalization.
Our model architecture is presented in Fig. 1, where the backbone is a vision transformer and parameterized by $\theta_{f}$, the task head predicts the AUs and the ID head outputs the subject identities, respectively. The input image $\boldsymbol{x}$ is first mapped by the backbone $G_{f}(\cdot;\theta_{f})$ to a D-dimensional feature vector $\boldsymbol{f}\in\mathbb{R}^{D}$, then the feature vector $\boldsymbol{f}$ is fed into the task head $G_{f}(\cdot;\theta_{au})$ and the ID head $G_{f}(\cdot;\theta_{id})$. simultaneously. Assume that we have data samples $(\boldsymbol{x},y,d)\sim D_{s}$, parameters $\theta_{au}$ of the task head are optimized to minimize the AU loss $L_{au}$ given AU label $y$, and parameters $\theta_{id}$ of the ID head are trained to minimize the identity loss $L_{id}$ given the identity label $d$.
To make the feature vector $\boldsymbol{f}$ invariant to subject identity, we seek the parameters $\theta_{f}$ of the backbone that maximize the identity loss $L_{id}$ (Eq. 3), so that the backbone excludes the identity-based features. In the meantime, the backbone $G_{f}(\cdot;\theta_{f})$ is expected to minimize the AU loss $L_{au}$. Formally, we consider the following functional loss:
@@
Figure 1: Architecture of Identity Adversarial Training. The task head and ID head both are a linear classifier predicting the AUs and identity, respectively. During training, the task head is optimized by $\frac{\partial L_{au}}{\partial\theta_{f}}$ and the ID head is optimized by $\frac{\partial L_{id}}{\partial\theta_{f}}$. The gradient reverse layer multiplies the gradient by a negative value $-\lambda$ to unlearn the features capable of recognizing identities. Finally, the parameters of the backbone are optimized by the two forces: $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$ and $\frac{\partial L_{au}}{\partial\theta_{f}}$.
@@
where the parameter $\lambda$ controls the trade-off between the two objectives that shape the feature $\boldsymbol{f}$ during learning. Comparing the identity loss $L_{id}$ in Eq. 3 and Eq. 4, $\theta_{f}$ is optimized to maximize to increase $L_{id}$ while $\theta_{id}$ is learned to reduce $L_{id}$. To achieve these two opposite optimizations through regular gradient descent and backpropagation, the gradient reverse layer is designed to reverse the identity partial derivative $\frac{\partial L_{id}}{\partial\theta_{id}}$ before it is propagated to the backbone. The resultant derivative $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$, together with $\frac{\partial L_{au}}{\partial\theta_{f}}$, are used to update the backbone parameter $\theta_{f}$.
@@
Intuitively, the backbone is still optimized to learn the AU-related features, but under the force of reducing the identity-related features. The ‘Min-Max’ training paradigm in gradient reverse (see Eq. 3) resembles the adversarial training [50] and Generative Adversarial Networks (GANs) [24], so we name our method ‘Identity Adversarial Training’ for the AU detection task.
@@
Importantly, we reveal the key design of identity adversarial training for AU detection: a strong adversarial regularization (large magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$) is required to learn identity-invariant features for the backbone. Specifically, we propose to use a large $\lambda$ and a linear projection layer for the ID head. The former scales up the $-\frac{\partial L_{id}}{\partial\theta_{f}}$ and the latter ensures a large $L_{id}$, leading to a large $||-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}||$ during training. In Sec. 5.3, we will show that the small $\lambda$ and 2-layer MLP ID head used by [87] would lead to a weak identity regularization (small magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$) and inferior AU performance. We defer more details and analysis to Sec. 5.3
4 Experiments
We test the performance of FMAE and FMAE-IAT on AU benchmarks, using the F1 score. To illustrate the generalization of FMAE, we also report its facial expression recognition (FER) accuracy on RAF-DB [60] and AffectNet [54] databases.
4.1 Datasets
4.2 Implementation details
4.3 Result of FMAE
We first show the F1 score of FMAE on the BP4D dataset in Tab. 2. FMAE achieves the same average F1 (66.6%) with the state-of-the-art method MDHR [70] which utilizes a two-stage model to learn the hierarchical AU relationships. Here, we see the effectiveness of data-centric facial representation learning, and demonstrate that a simple vision transformer [18], which is the architecture of FMAE, is capable of learning complex AU relationships. FMAE surpasses all previous work on BP4D+ and DISFA by achieving 66.2% and 68.7% F1 scores, respectively (see Tab. 3 and Tab. 4).
To further verify the importance of the Face9M dataset, we compare FMAE pretrained on Face9M with FMAE pretrained on ImageNet-1k [17], using BP4D as the test set. Figure 2 shows that FMAE pretrained on Face9M always outperforms the one pretrained on ImageNet-1k given the same model size (ViT-base or ViT-large). Also, we empirically demonstrate that FMAE benefits from the scaling effect of model size on AU detection tasks (see the green line in Fig. 2).
In addition to AU detection, we also benchmark FMAE on the downstream facial task of FER. We present the results of FMAE on AffectNet-8 and RAF-DB in Tab. 1. FMAE sets a new state-of-art accuracy on both datasets (65% on AffectNet-8 and 93.09% on RAF-DB). Note that, we did not test FMAE-IAT on FER tasks because these datasets do not include the identity labels and do not suffer from identity shortcut learning due to the large number of subjects.
Table 1: Results of accuracy on FER benchmarks. FMAE surpasses all previous work.
Methods Venue AffectNet-8 RAF-DB
DMUE [62] CVPR’21 62.84 88.76
EAC [84] ECCV’22 - 89.99
SOFT [46] ECCV’22 62.69 90.42
CLEF [80] ICCV’23 62.77 90.09
LA-Net [71] ICCV’23 64.54 91.78
FRA [22] CVPR’24 - 90.76
FMAE (ours) 65.00 93.09
@@ 4.3 Result of FMAE
4.4 Result of FMAE-IAT
Although FMAE has already achieved superior results on AU benchmarks, we highlight that the Identity Adversarial Training could further boost the performance of FMAE across all AU datasets. Specifically, we compare FMAE-IAT with the most recent state of the art methods on BP4D, BP4D+ and DISFA datasets. Tab. 2 suggests that FMAE-IAT shows superior performance by achieving an average F1 Score of 67.1% and FMAE-IAT ranks as the best or second-best performer in several individual AUs, notably AU 1, 2, 4, 12, 17 and 23. Similarly, FMAE-IAT also stands out on BP4D+ dataset with the highest average F1 score of 66.8%. Our results on the DISFA benchmark are even more distinguishing, FMAE-IAT gains the best or second-best performance on 6 out of 8 AUs, pushing the average F1 score beyond the 70% mark.
For the gradient reverse layer, we use $\lambda=2.0$ for BP4D, $\lambda=1.0$ for BP4D+ and $\lambda=0.5$ for DISFA. Differing from the setting $\lambda\in[0.008,0.08]$ used in [87], we emphasize that a strong IAT regularization is necessary for AU tasks and we defer the in-depth discussion throughout Sec. 5.

### v2

@@ Abstract
Facial Action Unit (AU) detection has gained significant attention as it enables the breakdown of complex facial expressions into individual muscle movements. In this paper, we revisit two fundamental factors in AU detection: diverse and large-scale data and subject identity regularization. Motivated by recent advances in foundation models, we highlight the importance of data and introduce Face9M, a diverse dataset comprising 9 million facial images from multiple public sources. Pretraining a masked autoencoder on Face9M yields strong performance in AU detection and facial expression tasks. More importantly, we emphasize that the Identity Adversarial Training (IAT) has not been well explored in AU tasks. To fill this gap, we first show that subject identity in AU datasets creates shortcut learning for the model and leads to sub-optimal solutions to AU predictions. Secondly, we demonstrate that strong IAT regularization is necessary to learn identity-invariant features. Finally, we elucidate the design space of IAT and empirically show that IAT circumvents the identity-based shortcut learning and results in a better solution. Our proposed methods, Facial Masked Autoencoder (FMAE) and IAT, are simple, generic and effective. Remarkably, the proposed FMAE-IAT approach achieves new state-of-the-art F1 scores on BP4D (67.1%), BP4D+ (66.8%), and DISFA (70.1%) databases, significantly outperforming previous work. We release the code and model at https://github.com/forever208/FMAE-IAT.
@@ 1 Introduction
The Facial Action Coding System (FACS) was developed to objectively encode facial behavior through specific movements of facial muscles, named Action Units (AU) [19]. Compared with facial expression recognition (FER) [37, 84, 35] and valence and arousal estimation [85, 51, 54], detecting action units offers a more nuanced and detailed understanding of human facial behavior capturing multiple individual facial actions simultaneously.
@@
This problem attracted considerable interest within the deep learning community [13, 32, 72, 56, 61, 73]. Many works used a facial region prior [39, 56, 13], introduced extra modalities [80, 69, 81], or incorporated the inherent AU relationships [38, 46, 75] to solve the AU detection task and achieved significant advancements. Diverging from these approaches, which often necessitate complex model designs or depend heavily on prior AU knowledge, in this paper, we revisit two fundamental factors that significantly contribute to the AU detection task: diverse and large-scale data and subject identity regularization.
@@
Recently, data has become pivotal in training foundation models [57, 2, 42, 64] and large language models [8, 1, 65, 15]. Following this trend, we introduce Face9M, a large-scale and diverse facial dataset curated and refined from publicly available datasets for pretraining. Different from contrastive learning methods [9, 13, 22], we propose to do facial representation learning using Masked Autoencoders (MAE) [27]. The underlying motivation is that most facial tasks require a fine-grained understanding of the face, and masked pretraining results in lower-level semantics than contrastive learning according to  [5]. Our large-scale facial representation learning approach demonstrates excellent generalization and scalability in downstream tasks. Notably, our proposed Facial Masked Autoencoder (FMAE), pretrained on Face9M, sets new state-of-the-art benchmarks in both AU detection and FER tasks.
@@
Similar to the importance of data, domain knowledge and task-prior knowledge can be incorporated into the model in the form of regularization [26, 31, 52, 55] to improve task performance. Our key observation is that popular AU detection benchmarks (i.e. BP4D [83], BP4D+ [86], DISFA [50]) include at most 140 human subjects and 192,000 images, meaning that each subject has hundreds of annotated images. This abundance can lead models to prefer simple, easily recognizable patterns over more complex but generalizable ones, as suggested by the shortcut learning theory [23, 29]. Therefore, we hypothesize that AU detection models tend to learn the subject identity features to infer the AUs, resulting in learning a trivial solution that does not generalize well. To verify our hypothesis, we employed the linear probing technique — adding a learnable linear layer to a trained AU model while freezing the network backbone —to measure identity recognition accuracy. The high accuracy (83%) we obtained in predicting the identities of the subjects clearly shows that the models effectively ‘memorize’ subject identities. To counteract the learning of identity-based features, we propose in this paper Identity Adversarial Training (IAT) for AU detection task by adding a linear identity prediction head and unlearning the identity feature using gradient reverse [20]. Further analysis shows that IAT significantly reduces the identity accuracy of linear probing and leads to better learning dynamics that avoid convergence to trivial solutions. This method further improves our AU models beyond the advantages brought by pretraining with a large-scale dataset.
@@
  • We highlight the identity shortcut learning issue and propose the use of a linear identity head and a large gradient reversal scalar in IAT to mitigate this issue for AU detection.
  • We release the code and checkpoint of FMAE with various model sizes (small, base, large), aiming at facilitating all facial tasks.
@@ 3.1 Large-scale Facial MAE Pretraining
While the machine learning community has long established the importance of having rich and diverse data for training, recent successes in foundation models and large language models illustrated the full potential of pretaining [57, 42, 8, 1, 65]. In line with this, our research pivots towards a nuanced exploration of data diversity and quantification in the context of facial representation learning. Unlike natural image datasets like ImageNet-1k, face datasets have low variance. Also, we observe that different facial datasets have domain shifts regarding the facial area, perspective and background. To increase the data diversity, we propose to collect a large facial dataset for pertaining from multiple data sources.
We first collect facial images from CelebA  [44], FFHQ  [33], VGGFace2 [11], CASIA-WebFace [76], MegaFace [34], EDFace-Celeb-1M [79], UMDFaces [6] and LAION-Face [89] datasets, because these datasets contain a massive number of identities collected in diverse scenarios. For instance, the facial images in UMDFaces also capture the upper body with various image sizes, while some datasets (FFHQ, CASIA-WebFace) mainly feature the center face. We then discard images whose width-height-ratio or height-width-ratio is larger than 1.5. Finally, the remaining images are resized to 224*224. The whole process yields 9 million facial images (termed Face9M) which will be used for self-supervised facial pertaining.
Regarding representation learning methods, we apply Masked Image Modeling (MIM) [27] as it tends to learn more fine-grained features than contrastive learning according to the study in [5], which benefits facial behavior understanding. Specifically, we utilize Face9M to train a masked autoencoder (MAE) by the mean squared error between the reconstructed and original images in the pixel space. The resulting model is termed FMAE, and the decoder of FMAE is discarded for the downstream tasks.
@@ 3.2 Identity Adversarial Training
One of our key findings in this paper is that, the limited number of subjects in AU datasets makes identity recognition a trivial task and provides a shortcut learning path, resulting in a AU model that contains identity-related features and does not generalize well (see Section 5). Motivated by the gradient reverse in domain adaption [20], we propose to apply gradient reverse on AU detection to learn identity-invariant features, aiming at better model generalization.
Our model architecture is presented in Figure 1, where the backbone is a vision transformer and parameterized by $\theta_{f}$, the AU head predicts the AUs and the ID head outputs the subject identities, respectively. The input image $\boldsymbol{x}$ is first mapped by the backbone $G_{f}(\cdot;\theta_{f})$ to a D-dimensional feature vector $\boldsymbol{f}\in\mathbb{R}^{D}$, then the feature vector $\boldsymbol{f}$ is fed into the AU head $G_{f}(\cdot;\theta_{au})$ and the ID head $G_{f}(\cdot;\theta_{id})$. simultaneously. Assume that we have data samples $(\boldsymbol{x},y,d)\sim D_{s}$, parameters $\theta_{au}$ of the AU head are optimized to minimize the AU loss $L_{au}$ given AU label $y$, and parameters $\theta_{id}$ of the ID head are trained to minimize the identity loss $L_{id}$ given the identity label $d$.
To make the feature vector $\boldsymbol{f}$ invariant to subject identity, we seek the parameters $\theta_{f}$ of the backbone that maximize the identity loss $L_{id}$ (Equation 3), so that the backbone excludes the identity-based features. In the meantime, the backbone $G_{f}(\cdot;\theta_{f})$ is expected to minimize the AU loss $L_{au}$. Formally, we consider the following functional loss:
@@
where the parameter $\lambda$ controls the trade-off between the two objectives that shape the feature $\boldsymbol{f}$ during learning. Comparing the identity loss $L_{id}$ in Equation 3 and Equation 4, $\theta_{f}$ is optimized to maximize to increase $L_{id}$ while $\theta_{id}$ is learned to reduce $L_{id}$. To achieve these two opposite optimizations through regular gradient descent and backpropagation, the gradient reverse layer is designed to reverse the identity partial derivative $\frac{\partial L_{id}}{\partial\theta_{id}}$ before it is propagated to the backbone. The resultant derivative $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$, together with $\frac{\partial L_{au}}{\partial\theta_{f}}$, are used to update the backbone parameter $\theta_{f}$.
@@
Intuitively, the backbone is still optimized to learn the AU-related features, but under the force of reducing the identity-related features. The ‘Min-Max’ training paradigm in gradient reverse (see Equation 3) resembles the adversarial training [49] and Generative Adversarial Networks (GANs) [24], so we name our method ‘Identity Adversarial Training’ for the AU detection task.
Figure 1: Architecture of Identity Adversarial Training. The AU head and ID head both are a linear classifier predicting the AUs and identity, respectively. The backbone $G_{f}(\cdot;\theta_{f})$ is the encoder of the pretrained FMAE. During training, the AU head is optimized by $\frac{\partial L_{au}}{\partial\theta_{f}}$ and the ID head is optimized by $\frac{\partial L_{id}}{\partial\theta_{f}}$. The gradient reverse layer multiplies the gradient by a negative value $-\lambda$ to unlearn the features capable of recognizing identities. Finally, the parameters of the backbone are optimized by the two forces: $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$ and $\frac{\partial L_{au}}{\partial\theta_{f}}$.
@@
Importantly, we reveal the key design of identity adversarial training for AU detection: a strong adversarial regularization (large magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$) is required to learn identity-invariant features for the backbone. Specifically, we propose to use a large $\lambda$ and a linear projection layer for the ID head. The former scales up the $\frac{\partial L_{id}}{\partial\theta_{f}}$ and the latter ensures a large $L_{id}$, leading to a large $||-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}||$ during training. In Section 5.3, we will show that the small $\lambda$ and 2-layer MLP ID head used by [87] would lead to a weak identity regularization (small magnitude of $-\lambda\frac{\partial L_{id}}{\partial\theta_{f}}$) and inferior AU performance. We defer more details and analysis to Section 5.3
@@
4 Experiments
We test the performance of FMAE and FMAE-IAT on AU benchmarks, using the F1 score. To illustrate the representation learning efficacy of FMAE, we also report its facial expression recognition (FER) accuracy on RAF-DB [59] and AffectNet [53] databases, and compare FMAE with previous face models pretrained based on contrastive learning.
4.1 Datasets
4.2 Implementation details
4.3 Result of FMAE
We first show the F1 score of FMAE on the BP4D dataset in Table 1. FMAE achieves the same average F1 (66.6%) with the state-of-the-art method MDHR [69] which utilizes a two-stage model to learn the hierarchical AU relationships. Here, we see the effectiveness of data-centric facial representation learning, and demonstrate that a simple vision transformer [18], which is the architecture of FMAE, is capable of learning complex AU relationships. FMAE surpasses all previous work on BP4D+ and DISFA by achieving 66.2% and 68.7% F1 scores, respectively (see Table 2 and Table 3).
@@ 4.3 Result of FMAE
To further verify the importance of the Face9M dataset, we compare FMAE pretrained on Face9M with FMAE pretrained on ImageNet-1k [17], using BP4D as the test set. Figure 2 shows that FMAE pretrained on Face9M always outperforms the one pretrained on ImageNet-1k given the same model size (ViT-base or ViT-large). Also, we empirically demonstrate that FMAE benefits from the scaling effect of model size on AU detection tasks (see the green line in Figure 2).
Table 4: Results of accuracy on FER benchmarks. FMAE surpasses all previous contrastive-related work.
Model Contrastive MIM AffectNet-8 RAF-DB
MCF [68] $\surd$ $\surd$ 60.98 86.86
FaRL [89] $\surd$ $\surd$ - 88.31
CLEF [80] $\surd$ 62.77 90.09
FRA [22] $\surd$ - 90.76
LA-Net [70] $\surd$ 64.54 91.78
FMAE (ours) $\surd$ 64.79 93.45
In addition to AU detection, we benchmark FMAE on the downstream facial task of FER to verify the effectiveness of masked image representation learning. We present the results of FMAE on AffectNet-8 and RAF-DB in Table 4 and compare FMAE with other contrastive learning-based models. FMAE sets a new state-of-art accuracy on both datasets (64.79% on AffectNet-8 and 93.45% on RAF-DB). Note that, we did not test FMAE-IAT on FER tasks because these datasets do not include the identity labels and do not suffer from identity shortcut learning due to the large number of subjects.
4.4 Results of FMAE-IAT
Although FMAE has already achieved superior results on AU benchmarks, we highlight that the Identity Adversarial Training could further boost the performance of FMAE across all AU datasets. Specifically, we compare FMAE-IAT with the most recent state of the art methods on BP4D, BP4D+ and DISFA datasets. Table 1 suggests that FMAE-IAT shows superior performance by achieving an average F1 Score of 67.1% and FMAE-IAT ranks as the best or second-best performer in several individual AUs, notably AU 1, 2, 4, 12, 17 and 23. Similarly, FMAE-IAT also stands out on BP4D+ dataset with the highest average F1 score of 66.8% shown in Table 2. Our results on the DISFA benchmark given in Table 3 are even more distinguishing, FMAE-IAT gains the best or the second-best performance on 6 out of 8 AUs, pushing the average F1 score beyond the 70% mark.
For the gradient reverse layer, we use $\lambda=2.0$ for BP4D, $\lambda=1.0$ for BP4D+ and $\lambda=0.5$ for DISFA. Differing from the setting $\lambda\in[0.008,0.08]$ used in [87], we emphasize that a strong IAT regularization is necessary for AU tasks and we defer the in-depth discussion throughout Section 5.